OpenStudy (anonymous):

a 3 digit number is formed from the odd digits 1,3,5,7,9. numbers can only be used once how many numbers are larger than 350

ganeshie8 (ganeshie8):

_ _ _ left most space can be chosen from {3, 5, 7, 9} = 4 ways next space can be chosen from {5, 7, 9}-1 = 2 ways last spce canbe chosen from {1, 3, 5, 7, 9}-2 = 3 ways total numbers larger than 350 = 4*2*3

OpenStudy (anonymous):

howcome answer is 4*3*3? what to do with the 1st and 2nd digit when theres overlap in set solutions??

ganeshie8 (ganeshie8):

wait.. i see something wrong ... gimme some time

ganeshie8 (ganeshie8):

pls scratch previous stuff.. see if below makes sense : case I : first digit = 3 second digit can be chosen from {5, 7, 9} = 3 ways last digit can be chosen from {1, 3, 5, 7, 9} -2 = 3 ways total number of ways = 3*3 = 9 case II : first digit can be one of { 5, 7, 9} = 3 ways second digit can be chosen from {5, 7, 9} -1 = 2 ways last digit can be chosen from {1, 3, 5, 7, 9} -2 = 3 ways total number of ways = 3*2*3 = 18 total numbers larger than 350 = 9 + 18

OpenStudy (anonymous):

yeah i tried that.. but the answer says 36

ganeshie8 (ganeshie8):

oh

ganeshie8 (ganeshie8):

lets see if some flaw there in our logic..

OpenStudy (anonymous):

and a question.. for the second case where first digit is from {5,7,9}, should the set for second digit be ({5, 7, 9} -1) or {1,3,5,7,9}-1 ? isnt any number with 5,7,9 in first digit larger than 350 so second digit can be any ?

ganeshie8 (ganeshie8):

awesome ! you're right !! but still we get answer as 36+9 = 45 which is incorrect.. .

OpenStudy (anonymous):

yeah i tried that.. not sure.

OpenStudy (anonymous):

its okay if you cant get it i'll ask teacher tomorrow

ganeshie8 (ganeshie8):

45 looks correct to me

ganeshie8 (ganeshie8):

ok :D

OpenStudy (anonymous):

and for these problems, do you do break the problem into separates cases or ?

ganeshie8 (ganeshie8):

yeah here we must break it into 2 cases. it cant get more simpler than this i think...

OpenStudy (anonymous):

choose the leftmost digit, now there are 12 possibilities for that digit. 5,7,9 each give 12 numbers > 350 3 gives (12-3) numbers > 350 --- we lose the three for 31- 1 gives 0 numbers > 350 I get 45. I think you are right @ganeshie8

ganeshie8 (ganeshie8):

yes. thanks @telliott99 just enumerated all the permutations. 45 is correct ! PFA

OpenStudy (anonymous):

thanks!! what does PFA mean btw

ganeshie8 (ganeshie8):

Please Find Attached when you attach something with ur reply you indicate that with PFA :D